How many ways can a group of 7 people be chosen from a class of 15?
Answer by jim_thompson5910(35256)

n C r = (n!)/(r!(n-r)!)

15 C 7 = (15!)/(7!*(15-7)!)

15 C 7 = (15!)/(7!*8!)

15 C 7 = (15*14*13*12*11*10*9*8!)/(7!*8!)

15 C 7 = (15*14*13*12*11*10*9)/(7!)

15 C 7 = (15*14*13*12*11*10*9)/(7*6*5*4*3*2*1)

15 C 7 = (32432400)/(5040)

15 C 7 = 6435


Answer: 6435